OpenStudy (anonymous):

Let g(x) be the reflection of f(x) = x2 +3 in the x-axis. What is the function rule for g(x)?

OpenStudy (anonymous):

A. g(x) = -x2 - 3 B. g(x) = -x2 + 3 C. g(x) = x2 - 3 D. g(x) = x2 + 3

OpenStudy (hyper):

A

OpenStudy (anonymous):

are u sure

OpenStudy (hyper):

is that an exponent for the x ?

OpenStudy (anonymous):

yes

OpenStudy (hyper):

k, then its A and this could be proved by tilting the curve of x^2 + 3 so that is is reflected by the x axis. In effect, it becomes -(x^2 +3) = A

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = 1/3x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7